what is the equation in standard form that passes through the point (2,5) and has the same slope as the graph of 3x+y=9?
Answer by mananth(16946)

3 x + 1 y = 9
Find the slope of this line
make y the subject
1 y = -3 x + 9
Divide by 1
y = -3 x + 9
Compare this equation with y=mx+b
slope m = -3

m= -3 ,point ( 2 , 5 )
Find b by plugging the values of m & the point in
y=mx+b
5 = -6 + b
b= 11
m= -3
Plug value of the slope and b in y = mx +b
The required equation is, y = -3 x + 11

m.ananth@hotmail.ca